As the new accountant for Cohen & Co., you have been asked to provide a succinct analysis of financial performance for the yeActual operating income 1 Master budget operating income 2. Total master budget variance 3 Total flexible-budget variance 4.
Answer #1
Actual Master (static) budget Flexible budget
Units sold 45,000 50,000 45,000
Sales 412,000 458,000 412,200 *
(-) variable cost (242,000) (278,000) (250,200)*
Contribution margin 170,000 180,000 162,000
(-) Fixed cost (151,000) (143,000) (143,000)*
Operating income 19,000 37,000 19,000

* 412,200= (458,000×45,000)÷50,000

* 250,200= (278,000×45,000)÷50,000

* Budgeted fixed cost does not change with output level.

Actual operating income 19,000
Master budget operating income 37,000
Total master budget variance 18,000 U
Total flexible-budget variance 0 none
Sales volume variance 18,000 U

Sales volume variance = Flexible budget operating income- Master budget operating income

= 19,000-37,000

=18,000 (U)

Total flexible-budget variance=Actual operating income- flexible budget operating income

=19,000-19,000

=0.

Total master budget variance= sales volume variance+ flexible budget variance

= (18,000)+(0)

= 18,000 (U)

____×____

All the best

Any doubt regarding the answer comment below